# CrackSegFlow: Controllable Flow Matching Synthesis for Generalizable Crack Segmentation with a 50K Image–Mask Benchmark

### v2

@@ Abstract
Automated crack segmentation is essential for condition assessment, yet deployment is limited by scarce pixel-level labels and domain shift. We present CrackSegFlow, a controllable Flow Matching synthesis framework that generates crack images conditioned on binary masks with mask–image alignment. The renderer combines topology-preserving mask injection with edge gating to maintain thin-structure continuity and suppress false positives. A class-conditional Flow Matching mask model synthesizes masks with control over crack coverage, enabling balanced, topology-diverse data without manual annotation. We inject masks into crack-free backgrounds to diversify illumination and reduce false positives. On five datasets with a CNN–Transformer backbone, incorporating synthesized pairs improves in-domain performance by 5.37 mIoU and 5.13 F1, and target-guided cross-domain synthesis yields gains of 13.12 mIoU and 14.82 F1 using target mask statistics. We also release CSF-50K, 50,000 image–mask pairs for benchmarking.
@@ 1 Introduction
Automated crack inspection is critical for infrastructure condition assessment and maintenance. Advances in computer vision offer accurate, cost-efficient alternatives to manual crack detection and segmentation. Traditional image-processing methods relying on hand-crafted filters or thresholding rules are sensitive to noise and illumination variations Yamaguchi and Hashimoto [2010], Fujita and Hamamoto [2011], while 3D reconstruction approaches can deliver detailed geometric measurements Jahanshahi and Masri [2012] but remain resource-intensive to collect and process. As a result, research has shifted toward deep learning (DL) methods that leverage strong feature extraction capabilities to classify Dais et al. [2021], detect Qiu and Lau [2023], Jiang et al. [2024a], Mayya and Alkayem [2025], and segment Wang et al. [2024b], Zhang et al. [2025], Goo et al. [2025], Fan et al. [2025] cracks. Most notably, pixel-level segmentation directly provides the geometry (length, width, connectivity) needed for severity metrics, eliminating additional post-processing by assigning each pixel as crack or background.
@@ 5 Conclusion
This paper introduced CrackSegFlow, a controllable Flow Matching framework that renders realistic pavement images conditioned on an input crack mask while maintaining tight mask–image alignment. CrackSegFlow consists of two rendering modules—a topology-preserving mask injection module and a boundary-gated rendering module—that jointly enforce structural fidelity (connectivity and thin-structure continuity) and suppress mask-boundary artifacts, enabling the generated image to accurately match the provided mask under diverse pavement appearances. Because CrackSegFlow is explicitly mask-conditioned, the conditioning mask can be either real or synthetic, providing a unified mechanism for controllable data synthesis.

### v3

@@ Abstract
Defect segmentation is central to computer vision–based inspection of infrastructure assets during both construction and operation. However, deployment remains limited due to scarce pixel-level labels and domain shift across environments. We introduce CrackSegFlow, a controllable Flow Matching synthesis method that renders synthetic images of cracks from masks with pixel-level alignment. Our renderer combines topology-preserving mask injection with edge gating to maintain thin-structure continuity. Class-conditional FM samples masks for topology diversity, and CrackSegFlow renders aligned ground truth images from them. We further inject cracks onto crack-free backgrounds to diversify confounders and reduce false positives. Across five datasets and using a CNN–Transformer backbone, our results demonstrate that adding synthesized pairs improves in-domain performance by +5.37 mIoU / +5.13 F1, while target-guided cross-domain synthesis—driven by target mask statistics—adds +13.12 mIoU / +14.82 F1. We also release CSF-50K, a benchmark dataset, comprising 50,000 image–mask pairs.
@@ 1 Introduction
Automated crack inspection is critical for infrastructure condition assessment and maintenance, and it is increasingly used in automated visual inspection workflows (e.g., UAV/robot/vehicle imaging) to support maintenance prioritization and asset-management decisions. Advances in computer vision offer accurate, cost-efficient alternatives to manual crack detection and segmentation. Traditional image-processing methods relying on hand-crafted filters or thresholding rules are sensitive to noise and illumination variations Yamaguchi and Hashimoto [2010], Fujita and Hamamoto [2011], while 3D reconstruction approaches can deliver detailed geometric measurements Jahanshahi and Masri [2012] but remain resource-intensive to collect and process. As a result, research has shifted toward deep learning (DL) methods that leverage strong feature extraction capabilities to classify Dais et al. [2021], detect Qiu and Lau [2023], Jiang et al. [2024a], Mayya and Alkayem [2025], and segment Wang et al. [2024b], Zhang et al. [2025], Goo et al. [2025], Fan et al. [2025] cracks. Most notably, pixel-level segmentation directly provides the geometry (length, width, connectivity) needed for severity metrics, eliminating additional post-processing by assigning each pixel as crack or background. In Automation in Construction settings, such dense outputs enable consistent, scalable condition assessment and inform maintenance prioritization and asset-management decisions. Therefore, reducing pixel-level labeling cost and improving robustness under domain shift across sites, sensors, and surface textures are key requirements for deployable crack inspection systems.
@@ 5 Conclusion
Crack segmentation is a foundational step in automated visual inspection pipelines (e.g., UAV/robot/vehicle imaging) used for condition assessment of pavements and other civil infrastructure, and it directly supports downstream maintenance prioritization and asset-management decisions. By reducing pixel-level labeling burden and improving cross-domain robustness through controllable, mask-aligned synthesis, the proposed framework strengthens the reliability and scalability of automated inspection systems deployed across sites, sensors, and surface textures. This paper introduced CrackSegFlow, a controllable Flow Matching framework that renders realistic pavement images conditioned on an input crack mask while maintaining tight mask–image alignment. CrackSegFlow consists of two rendering modules—a topology-preserving mask injection module and a boundary-gated rendering module—that jointly enforce structural fidelity (connectivity and thin-structure continuity) and suppress mask-boundary artifacts, enabling the generated image to accurately match the provided mask under diverse pavement appearances. Because CrackSegFlow is explicitly mask-conditioned, the conditioning mask can be either real or synthetic, providing a unified mechanism for controllable data synthesis.